Opened 12 years ago

Closed 10 years ago

#148 closed flight processing (fixed)

UR08/01, flight day 140, Black_Venn

Reported by: mark1 Owned by: mark1
Priority: immediate Milestone: 2008 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/in_progress/2008/flight_data/uk/UR08_01-2008_140_Black_Venn

Data arrived from ARSF via SATA disk transfer 28th May 2008. (Initially downloaded from ARSF Friday 23rd May 2008)

PI: Stuart Marsh

Urgent processing request.

Sensors:

  • ATM (delivered 17th June 2008)
  • Eagle (delivered 17th June 2008)
  • Hawk (delivered 17th June 2008)

Quicklook

TBD.

Change History (7)

comment:1 Changed 12 years ago by mark1

Have a problem with the GPS Rinex data. It appears to end at 09:59:59 which is in the middle of acquisition period. Emailed Gary in hope of getting Rinex data for the period after 09:59:59

comment:2 Changed 12 years ago by mark1

Rinex data has now been supplied

comment:3 Changed 12 years ago by mark1

Hawk line 1 and Eagle line 2 (Data files SWIR..-2.hdr and VNIR..-3.hdr have NANs as GPS stop time. These have been replaced by approx end time from logsheet, and appears to work ok.

comment:4 Changed 11 years ago by mark1

  • Owner changed from mark1 to anee

Delivery directory created and located at

~mark1/scratch_space/UR08_01-2008_140_Black_Venn/

comment:5 Changed 11 years ago by anee

  • Owner changed from anee to mark1

Delivery checked, seems fine - rsynced to:

~/arsf_data/in_progress/2008/flight_data/uk/UR08_01-2008_140_Black_Venn

comment:6 Changed 11 years ago by mark1

  • Description modified (diff)

Data delivered on 2 DVDs 17th June 2008

comment:7 Changed 10 years ago by mggr

  • Resolution set to fixed
  • Status changed from new to closed

Data sent to NEODC for archiving.

Note: See TracTickets for help on using tickets.